A class has 12 boys and 21 girls. The class as a whole has a GPA (grade point average) of 2.96, and the boys have a GPA of 2.40. What is the GPA of the girls?

Answer by fractalier(6550)

This is a weighted average problem...we set it up like this...
(12+21)(2.96) = 12(2.40) + 21x
where x is the GPA of the girls...
33(2.96) = 28.8 + 21x
97.68 = 28.8 + 21x
68.88 = 21x
x = 68.88/21 = 3.28 (the girls' GPA)